OpenStudy (anonymous):

need help with simplifying 5x+4/x2-64+3/x-8

OpenStudy (anonymous):

so this problem reads (5x+4/x^2) - (64+3/x-8)

OpenStudy (lgbasallote):

\(\large \frac{5x + 4}{x^2 - 64} + \frac{3}{x-8} = \frac{5x+4}{(x+8)(x-8)} + \frac{3}{x-8}\) based from that..can you identify the LCD? remember that the LCD must contain BOTH denominators of both fractions

OpenStudy (anonymous):

(5x+4/x^2-64)+(3/x-8)

OpenStudy (anonymous):

so the LCD is x-8

OpenStudy (anonymous):

Well, you can simplify x^2-64. Do you have an idea how?

OpenStudy (lgbasallote):

like i said the LCD must contain BOTH denominators...since x - 8 does NOT contain x + 8 which is part of the first denominator.. x - 8 is NOT the LCD

OpenStudy (anonymous):

you can make that 5x+4/(x-8)(x+8)

OpenStudy (anonymous):

Yep, so your problem now looks like [5x+4/(x-8)(x+8) + (3/x-8)] do you know how to cancel out things that are similar on top and bottom?

OpenStudy (anonymous):

yeah but you can't cancel anything out there

OpenStudy (anonymous):

Good. So what is missing that the denominatours aren't the same?

OpenStudy (anonymous):

x+8 isn't in the 2nd one

OpenStudy (anonymous):

So you must multiply the numorator and denominatour by x+8 so they are even

OpenStudy (anonymous):

how would that read?

OpenStudy (anonymous):

that would read 3(x+8)/(x-8)(x+8)

OpenStudy (anonymous):

Then you would simplify the top portion

OpenStudy (anonymous):

okay, so then you just cancel the demoninators?

OpenStudy (anonymous):

first simplify 3(x+8)

OpenStudy (anonymous):

3x+24

OpenStudy (anonymous):

yes. So now your bottoms are even, you can ignore those. just look at the numerators. What does your problem look thus far?

OpenStudy (anonymous):

would the final answer be 4(2x+7)?

OpenStudy (anonymous):

5x+4+3x+24 is what your next step would be. If you simplified those, you would get 8x+28. take the 4 so it would be 4(2x+7), so correct!

OpenStudy (anonymous):

yes! new favorite web site

OpenStudy (anonymous):

I know! I was so amazed when I found this place :D You understand it, right? (the problem, that is)

OpenStudy (anonymous):

yeah that really helped now i just have to post 1 more queestion lol

OpenStudy (anonymous):

lol maybe I'll help you on there too, but I think your good to close this question!